OpenStudy (loser66):

find real and imaginary parts of (1-i)^i Please, help

rvc (rvc):

consider Z=(1-i)^i take log

OpenStudy (loser66):

@SithsAndGiggles if we do that way, we don't have imaginary part, right? @rvc I did and got stuck

OpenStudy (anonymous):

Hold on, those rules don't hold up....

OpenStudy (loser66):

we know that \(z^a = e^{alogz}\)

OpenStudy (loser66):

\((1-i)^i = e^{ilog(1-i)}\)

OpenStudy (loser66):

and log (1-i) = ln|1-i| + i arg (-pi/4)

rvc (rvc):

\[\rm LogZ=i(\log(1-i))\\~~~~~=i(1/2\log2+i\pi/4)\]

myininaya (myininaya):

have you tried writing 1-i in polar form ? r*e^(i theta) form is what I'm talking about

rvc (rvc):

oops its negative pi/4

rvc (rvc):

THen multiply the i

OpenStudy (loser66):

but ln|1-i| = ln (sqrt 2), right?

OpenStudy (loser66):

then, \(e^{ilog(1-i)}= e^{i (ln\sqrt 2+i(-\pi/4))}\) expand it, I got a complicated answer :(

rvc (rvc):

\[\rm ~Z=(1-i)^i \\ logZ=i(\log[1-i]) \\ logZ=i(\log \sqrt2-i\pi/4) \\logZ=\pi/4+ilog \sqrt2\\Z=e^{\pi/4} \cdot e^{ilog \sqrt2} \]

rvc (rvc):

now write e^ix = cos(x)+isin(x)

OpenStudy (loser66):

I want the form z = 1-i because I still have part b, c, d, e to work on that form. If I do your way , z = (1-i)^i, I can't work further :(

rvc (rvc):

\[\log(a+ib)=\log(r)+i(\theta) \\ where~r= \sqrt{a^2+b^2}~\ \theta= \tan^{-1}(b/a)\]

OpenStudy (kainui):

I would just plug the number into this and simplify: \[\Re\{z\} = \frac{z+z^*}{2}\]\[\Im\{z\} = \frac{z-z^*}{2i}\] In case you think I'm lying to you, here are some things you can do to prove it to yourself now that you wouldn't need to do ever again once you know that above. ---- You can tell that the real part is real because: \[\Re\{z\}^* = \frac{(z+z^*)^*}{2} = \frac{z+z^*}{2} = \Re\{z\}\] Similarly purely imaginary numbers obey this, \[\Im\{z\}^* = - \Im\{z\}\] And check to make sure that \[z = \Re\{z\} + i \Im\{z\}\]

rvc (rvc):

whats the question?

OpenStudy (loser66):

This is the whole question b) find the domain of analicity of z^n and show that (z^a)'= az^(a-1) where z^(a-1) is the principal branch of z^(a-1) c) z^az^b = z^(a+b) ? d) \(z_1^az_2^a = (z_1z_2)^a? e) find the Taylor series for f(z) = z ^a at z =1. Determine the radius of convergence

rvc (rvc):

hmmm i have no idea about that :/

rvc (rvc):

@ganeshie8 please help :)

OpenStudy (loser66):

Thank you @rvc

rvc (rvc):

just a sec please im asking others to help :)

ganeshie8 (ganeshie8):

\(c\) is correct but \(d\) may not hold always

OpenStudy (loser66):

c is correct iff the real of z >0 since log z is undefined on negative value of x axis.

OpenStudy (loser66):

c is false since \(z^a = e^{alogz}\\z^b = e^{blogz}\\z^az^b = e^{(a+b) logz} \neq e^{(a+b)}\)

rvc (rvc):

z^a is the answer after separating the imaginary part n real?

rvc (rvc):

@Loser66 ?

OpenStudy (loser66):

what do you mean?

OpenStudy (loser66):

Actually, I got most of them. Now I am working on radius of convergence of Taylor series now. :)

rvc (rvc):

ohok